# Spatio-Temporal Graph Neural Network for Urban Spaces: Interpolating Citywide Traffic Volume

### v1

@@ Abstract
Reliable street-level traffic volume data, covering multiple modes of transportation, helps urban planning by informing decisions on infrastructure improvements, traffic management, and public transportation. Yet, traffic sensors measuring traffic volume are typically scarcely located, due to their high deployment and maintenance costs. To address this, interpolation methods can estimate traffic volumes at unobserved locations using available data. Graph Neural Networks have shown strong performance in traffic volume forecasting, particularly on highways and major arterial networks. Applying them to urban settings, however, presents unique challenges: urban networks exhibit greater structural diversity, traffic volumes are highly overdispersed with many zeros, the best way to account for spatial dependencies remains unclear, and sensor coverage is often very sparse. We introduce the Graph Neural Network for Urban Interpolation (GNNUI), a novel urban traffic volume estimation approach. GNNUI employs a masking algorithm to learn interpolation, integrates node features to capture functional roles, and uses a loss function tailored to zero-inflated traffic distributions. In addition to the model, we introduce two new open, large-scale urban traffic volume benchmarks, covering different transportation modes: Strava cycling data from Berlin and New York City taxi data. GNNUI outperforms recent, some graph-based, interpolation methods across metrics (MAE, RMSE, true-zero rate, Kullback-Leibler divergence) and remains robust from 90% to 1% sensor coverage. On Strava, for instance, MAE rises only from 7.1 to 10.5, on Taxi from 23.0 to 40.4, demonstrating strong performance under extreme data scarcity, common in real-world urban settings. We also examine how graph connectivity choices influence model accuracy.
@@ 1 Introduction
Reliable data on citywide traffic volumes is critical for transportation and urban planning decisions. Urban traffic data across modes can provide insights for targeted infrastructure improvements, traffic management, or improving the efficiency of public transportation (Leduc, 2008; Zheng et al., 2014). Traffic volume data is collected by sparsely distributed traffic sensors that count the number of vehicles passing by. Extending this monitoring to all street segments (a street section between two intersections) is largely infeasible due to the high costs associated with deploying and maintaining such devices. We explore how traffic volumes in locations without sensors can be interpolated employing Graph Neural Network s (GNN s).
@@
(a) Street network in Manhattan, NYC, comprising 8,156 street segments. The 21 traffic sensors are marked with dots.
(b) Distribution of taxi trip volume per street segment in Manhattan, NYC.
(c) Cycling network in Berlin, comprising 4,958 street segments. The 40 bicycle sensors are marked with dots.
(d) Distribution of Strava cycling volume per street segment in Berlin.
Figure 1: Urban street networks (left) and corresponding distributions of street-level traffic volume (right). It is important to note that the traffic and cyclist sensors count all passing motorized vehicles and cyclists, respectively—not just taxis or Strava users. Taxi and Strava data represent only subsets of total traffic. Nonetheless, we include the sensor locations to highlight the extent of sensor sparsity in urban settings.
@@
Previous work has shown that interpolation with GNN s performs well for highways and similar networks (Yao and Huang, 2023; Zheng et al., 2023; Wu et al., 2021a, b; Appleby et al., 2020). However, these networks differ profoundly from urban street networks, and correspondingly, the existing models overlook several critical aspects of urban contexts. Urban street networks exhibit greater infrastructural diversity than highways, making it essential to incorporate additional features that capture these differences. For example, the number of shops or the speed limit can influence traffic volumes. Moreover, the urban traffic data can be highly overdispersed, as in the case of specific transportation services and modes (e.g., cycling), with a significant proportion of zero values (Figure 1). Most existing models rely on the assumption of normally distributed data and adopt corresponding loss functions, which may not work for urban data. Furthermore, the higher density and structural complexity of urban networks complicate spatial dependency modeling. In poly-centric urban forms, for example, similar traffic patterns can emerge even between distant areas. Since GNN s are highly sensitive to graph construction, a systematic comparison of graph configurations is needed. Lastly, urban networks tend to exhibit far lower sensor density relative to the number of street segments when compared to highways. For instance, Manhattan’s street network comprises 8,156 segments, yet only 21 sensors recorded motorized traffic in 2023 (NYC Open Data, 2024). Berlin’s cycling network spans 4,958 segments, but cyclist volume was collected by just 22 permanent and 18 temporary sensors (Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin, 2024b) (see Figure 1). This is in stark contrast to highway networks, such as California’s highway network, commonly used through the PeMS benchmark dataset, with over 40,000 sensors (State of California, ).
@@
To address these challenges, we propose Graph Neural Network for Urban Interpolation (GNNUI), a model that interpolates traffic volumes while accounting for urban-specific characteristics. It captures spatial dependencies through graph structure. To model temporal dependencies and enable effective interpolation, we adopt a masking algorithm and time window approach from the IGNNK model by Wu et al. (2021a). To better capture the heterogeneity of urban environments, GNNUI incorporates diverse street-level features, such as infrastructure, points of interest, and weather. To more accurately model urban traffic distributions, we introduce a zero-inflated negative binomial distribution (Zhuang et al., 2022). Since interpolation targets are known in advance, the full graph, including nodes with missing values, is included during training, allowing the model to leverage their features and context from the outset. We share the full implementation of our approach in the accompanying repository for transparency and reproducibility https://github.com/silkekaiser/GNNUI.git, and to support it, we assemble and publish two large-scale, street-level urban traffic datasets: Strava cycling data from Berlin and New York City taxi data, each enriched with auxiliary urban features and available for download (Kaiser, 2025). In addition, we empirically address two further questions: What is the most effective graph representation for interpolation? and How well can traffic volumes be predicted under varying data scarcity? To answer the first, we compare different adjacency matrix configurations. For the second, we evaluate model performance across a range of sensor coverage levels, from sparse to near-complete. Most existing research focuses on settings with high sensor availability; however, real-world urban networks are characterized by extreme data sparsity, making performance under low coverage particularly important.
@@
  • We demonstrate that GNNUI outperforms baseline methods. We systematically evaluate its robustness across varying levels of sensor coverage and demonstrate that it maintains strong performance, especially under extreme data scarcity. This is critical for real-world urban settings where sensor infrastructure is typically sparse, and reliable estimation is most needed.
@@ 2 Related Work
Despite the strong focus on forecasting, GNNs also have potential for interpolation tasks. Spatial interpolation (also known as spatial kriging or out-of-sample imputation) aims to estimate traffic conditions in areas without direct sensor coverage. Yet, these interpolation tasks received considerably less attention than forecasting in the literature. Traditional interpolation methods include kriging methods (Krige, 1951; Cressie, 2015), a pioneering approach using spatial correlation between data points and and their observed labels. It assumes a Gaussian distribution, which is often not the case in real-world data. Although there are workarounds (Saito and Goovaerts, 2000), circumventing this assumption is not always feasible. Other approaches harvested linear spatial dependencies to interpolate unobserved nodes, such as K nearest neighbor (KNN) (Cover and Hart, 1967) or inverse distance weighting (IDW) (Lu and Wong, 2008). Lastly, another approach was matrix/tensor completion (Bahadori et al., 2014; Deng et al., 2016; Zhou et al., 2012). Crucially, often such methods are transductive, meaning they cannot predict new nodes that have not been seen before; instead, one must retrain a model for each new graph structure.
@@ 3.1 Model
Network Design. We model our network such that each street segment is thought to be equipped with a traffic volume sensor measuring the number of vehicles passing by (target variable). We use “street segment” interchangeably with “sensor”, under the simplifying assumption of a one-to-one mapping. In practice, a street segment may have multiple sensors; however, since our framework treats all sensors on the same segment as recording the same traffic volume, extending our approach to handle multiple sensors per segment is straightforward. The nodes in our graph represent the $n$ sensor locations in the training data (Figure 3). The edges and their weights are determined by the adjacency matrix $W\in\mathbb{R}^{n\times n}$. The adjacency matrix can represent distance or similarity, resulting in a fully connected graph (lower left of the figure), or be binary, with edges only between direct neighbors (lower right).
@@
Prediction Task. The goal of our model is to interpolate the traffic volume for street segments that lack sensor data. For this, the model is provided with a) the target variables and node features of seen (i.e., sensor-equipped), and b) the node features of the unseen (i.e., not equipped with sensors) segments. All predictions are made over a fixed time window. The model predicts the target variable for unseen and seen nodes across the time window. Wu et al. (2021a) compute the loss across all nodes, aiming to generalize the message-passing mechanism. Instead, we want our model to specialize in interpolating new locations. Thus, we compute the loss and the error metrics only on the interpolated nodes, while accounting for missing data (mathematical definition in Appendix A). For example, in Figure 2(a), the model uses the node features and the target variable of segments 1 to 7 (seen), to predict the target variable for segments 8 and 9 (unseen) on which the error metrics are computed.
Node Features and Target Variables. While most interpolation models rely primarily on the target variable from available sensors, we also incorporate node-level features. Features are important in urban contexts, where local characteristics significantly impact traffic volume. For example, a busy shopping street and a quiet residential street may be adjacent but have very different traffic volumes, differences that the model can only learn if features are included. We incorporate $k$ node features in our model. The data is available across $p$ time steps, the tensor $X\in\mathbb{R}^{n\times p\times k}$ contains the training node feature inputs for all sensors locations, while the target variable is stored in the tensor $T\in\mathbb{R}^{n\times p\times 1}$.
Time Window Approach to Model Temporal Dependency. To capture temporal dependencies, we adopt the fixed-length time window strategy from Wu et al. (2021a). For a given start time $t$, the model interpolates traffic volumes for the period $[t,t+h)={t,t+1,\dots,t+h-1}$, where $h$ is the time window size. This assumes temporal correlation of traffic patterns within each window.
@@
Training Through Masking Strategy.
To train the model for interpolation, we adopt a masking strategy inspired by Wu et al. (2021a). We briefly summarize it here and provide full and technical details in Appendix B. Missing values are common in real-world sensor data and can be caused by malfunctions, maintenance, or construction activities. For instance, after removing outliers and accounting for gaps, 1.3% of the Strava data and 1.0% of the taxi data are missing. The masking strategy handles the missing data by replacing it with zero. During training, we simulate the interpolation process by constructing multiple training samples. A sample is a small training instance that includes a) a selected time window of length $h$, and b) a subset of nodes, randomly divided into supposedly seen and unseen nodes. Each sample mimics a localized spatio-temporal snapshot of the network, allowing the model to repeatedly learn the interpolation task from different perspectives. To mimic the real prediction setting, we mask the target variables of the supposedly unseen nodes by setting them to zero, just as we represent missing values in the test phase. The model is trained to predict these masked values using the target variables and node features of the supposedly seen and the node features of the supposedly seen nodes. Loss is computed only on the masked nodes. Also different from Wu et al. (2021a), we ensure that time windows during training do not overlap, to make use of all training data and adopt a scheduled learning rate to improve convergence and stability.
Setting target variables to zero can be problematic because the model cannot distinguish between genuinely zero values, those set to zero due to masking, and those set to zero because they are missing. To resolve this ambiguity, and differing from the original masking strategy in Wu et al. (2021a), we append two additional binary indicators to the feature vector of each node. These indicators reflect the masking and missingness status, respectively. As a result, each node’s feature vector includes two additional entries, expanding $k$ by two.
@@
Test Data from Same Time as Training Data. The test data comes from the same time period as the training data (in both Figure 2(a) and 2(b)). We train the model on time steps $1$ through $p$. The test is then performed on segments 8 and 9 during the time period $1$ through $p$. Since the test time overlaps with training, the model has already seen parts of the input (i.e., $X_{t}^{s}$, $X_{t}^{u}$, and $T_{t}^{s}$ under consideration of truly missing values). While this setup focuses on spatial interpolation, it requires retraining the model if predictions are needed for a different time. This is particularly suited for retrospective urban planning applications, where historical traffic distributions inform infrastructure decisions. In contrast, an alternative setup could consist of the model generalizing to both unseen locations and future times, meaning that no test inputs are observed during training. This is valuable in applications requiring continuously updated traffic forecasts, such as dynamic traffic control. This is explained and parallel to the results in this study, evaluated in an ablation study in Appendix C.
@@ 3.2 Adjacency matrices
Graph Neural Networks work with graph-structured data and, therefore, greatly depend on how the graph is defined (Ye et al., 2022). We want to evaluate which adjacency matrix is most suitable for urban traffic interpolation. In this work, we investigate three types of adjacency matrices for urban traffic interpolation: binary, distance-based, and similarity-based. We chose them due to their simplicity, interpretability, and widespread use in the traffic modeling literature.
Binary adjacency matrices capture the local topology of the urban street network by connecting nodes (i.e., street segments) that are directly linked in space. They are widely used in the traffic domain due to their simplicity and ability to preserve physical road connectivity (Guo et al., 2019; Ye et al., 2022; Zhao et al., 2020). A binary adjacency matrix is defined using street segments $v_{i}$ and $v_{r}$ as:
$w_{ir}=\begin{cases}1,&\text{if }v_{i}\text{ and }v_{r}\text{ meet at an intersection},\\ 0,&\text{otherwise}.\end{cases}$ (1)
@@
$w_{ij}=\cos{(X_{i}^{infra},X_{j}^{infra})}=\frac{X_{i}^{infra}\cdot X_{j}^{infra}}{||X_{i}^{infra}||\cdot||X_{j}^{infra}||}$ (3)

### v2

@@ Abstract
Graph Neural Networks have shown strong performance in traffic volume forecasting, particularly on highways and major arterial networks. Applying them to urban settings, however, presents unique challenges: urban networks exhibit greater structural diversity, traffic volumes are highly overdispersed with many zeros, the best way to account for spatial dependencies remains unclear, and sensor coverage is often very sparse. We introduce the Graph Neural Network for Urban Interpolation (GNNUI), a novel urban traffic volume estimation approach. GNNUI employs a masking algorithm to learn interpolation, integrates node features to capture functional roles, and uses a loss function tailored to zero-inflated traffic distributions. In addition to the model, we introduce two new open, large-scale urban traffic volume benchmarks, covering different transportation modes: Strava cycling data from Berlin and New York City taxi data. GNNUI outperforms recent, some graph-based, interpolation methods across metrics (MAE, RMSE, true-zero rate, Kullback-Leibler divergence) and remains robust from 90% to 1% sensor coverage. For example, on the Strava dataset, the MAE increases only from 7.1 to 10.5, and on the Taxi dataset, from 23.0 to 40.4. These results demonstrate that GNNUI maintains strong performance despite extreme data scarcity, a common condition in real-world urban settings. We also examine how graph connectivity choices influence model accuracy.
@@ 1 Introduction
Reliable data on citywide traffic volumes is critical for transportation and urban planning decisions. Urban traffic data across modes can provide insights for targeted infrastructure improvements, traffic management, or improving the efficiency of public transportation (Leduc, 2008; Zheng et al., 2014). Traffic volume data is collected by sparsely distributed physical traffic sensors that count the number of vehicles passing by. Extending this monitoring to all street segments (a street section between two intersections) is largely infeasible due to the high costs associated with deploying and maintaining such devices. We explore how traffic volumes in locations without sensors can be interpolated employing Graph Neural Network s (GNN s), using datasets with full spatial and temporal coverage, taxi trip and Strava cycling counts, as proxies for overall traffic activity, given the limited availability of comprehensive ground-truth data.
@@
Previous work has shown that interpolation with GNN s performs well for highways and similar networks (Yao and Huang, 2023; Zheng et al., 2023; Wu et al., 2021a, b; Appleby et al., 2020). However, these networks differ profoundly from urban street networks, and correspondingly, the existing models overlook several critical aspects of urban contexts. Urban street networks exhibit greater infrastructural diversity than highways, making it essential to incorporate additional features that capture these differences. For example, the number of shops or the speed limit can influence traffic volumes. Moreover, the urban traffic data can be highly overdispersed, as in the case of specific transportation services and modes (e.g., cycling), with a significant proportion of zero values (Figure 1). Most existing models rely on the assumption of normally distributed data and adopt corresponding loss functions, which may not work for urban data. Furthermore, the higher density and structural complexity of urban networks complicate spatial dependency modeling. In poly-centric urban forms, for example, similar traffic patterns can emerge even between distant areas. Since GNN s are highly sensitive to graph construction, a systematic comparison of graph configurations is needed. Lastly, urban networks tend to exhibit far lower sensor density relative to the number of street segments when compared to highways. For instance, Manhattan’s street network comprises 8,156 segments, yet only 21 sensors recorded motorized traffic in 2023 (NYC Open Data, 2024). Berlin’s cycling network spans 4,958 segments, but cyclist volume was collected by just 22 permanent and 19 temporary sensors (Senate Department for the Environment, Mobility, Consumer and Climate Protection Berlin, 2024b) (see Figure 1). This is in stark contrast to highway networks, such as California’s highway network, commonly used through the PeMS benchmark dataset, with over 40,000 sensors (State of California, ).
@@
To address these challenges, we propose Graph Neural Network for Urban Interpolation (GNNUI), a model that interpolates traffic volumes while accounting for urban-specific characteristics. It captures spatial dependencies through graph structure. To model temporal dependencies and enable effective interpolation, we adopt a masking algorithm and time window approach from the IGNNK model by Wu et al. (2021a). To better capture the heterogeneity of urban environments, GNNUI incorporates diverse street-level features, such as infrastructure, points of interest, and weather. To more accurately model urban traffic distributions, we introduce a zero-inflated negative binomial distribution (Zhuang et al., 2022). Since interpolation targets are known in advance, the full graph, including nodes with missing values, is included during training, allowing the model to leverage their features and context from the outset. We share the full implementation of our approach in the accompanying repository for transparency and reproducibility https://github.com/silkekaiser/GNNUI.git, and to support it, we assemble and publish two large-scale, street-level urban traffic datasets: Strava cycling data from Berlin and New York City taxi data, each enriched with auxiliary urban features, offering complete spatio-temporal coverage across all street segments, and available for download (Kaiser, 2025). In addition, we empirically address two further questions: What is the most effective graph representation for interpolation? and How well can traffic volumes be predicted under varying data scarcity? To answer the first, we compare different adjacency matrix configurations. For the second, we evaluate model performance across a range of simulated sensor coverage levels, from sparse to near-complete. Most existing research focuses on settings with high sensor availability; however, real-world urban networks are characterized by extreme data sparsity, making performance under low coverage particularly important.
(a) Street network of Manhattan, New York City, comprising 8,156 street segments. The 21 existing traffic sensors are shown as dots. The inset in the upper right displays the distribution of taxi trip volumes per street segment.
(b) Cycling network of Berlin, comprising 4,958 street segments. The 41 existing bicycle sensors are shown as dots. The inset in the upper right displays the distribution of Strava cycling volumes per street segment.
Figure 1: Urban street networks (left) and corresponding distributions of street-level traffic volume (right). Street segments used for training and validation are shown in one color, and test segments in another, illustrating the spatial split applied in our experiments. Dots indicate the locations of the currently employed traffic and cycling sensors, which are displayed for reference only. Our analyses do not rely on these sensors; instead, we use trajectory-based data from taxi trips (New York) and Strava cycling records (Berlin). The sensors are shown solely to illustrate the sparsity of conventional urban monitoring infrastructure.
@@
  • We demonstrate that GNNUI outperforms baseline methods. We systematically evaluate its robustness across varying levels of simulated sensor coverage and demonstrate that it maintains strong performance, especially under extreme data scarcity. This is critical for real-world urban settings where sensor infrastructure is typically sparse, and reliable estimation is most needed.
@@ 2 Related Work
Despite the strong focus on forecasting, GNNs also have potential for interpolation tasks. Spatial interpolation (also known as spatial kriging or out-of-sample imputation) aims to estimate traffic conditions in areas without direct sensor coverage. Yet, these interpolation tasks received considerably less attention than forecasting in the literature. Traditional interpolation methods include kriging methods (Krige, 1951; Cressie, 2015), a pioneering approach using spatial correlation between data points and their observed labels. It assumes a Gaussian distribution, which is often not the case in real-world data. Although there are workarounds (Saito and Goovaerts, 2000), circumventing this assumption is not always feasible. Other approaches harvested linear spatial dependencies to interpolate unobserved nodes, such as K nearest neighbor (KNN) (Cover and Hart, 1967) or inverse distance weighting (IDW) (Lu and Wong, 2008). Another approach was matrix/tensor completion (Bahadori et al., 2014; Deng et al., 2016; Zhou et al., 2012). Crucially, often such methods are transductive, meaning they cannot predict new nodes that have not been seen before; instead, one must retrain a model for each new graph structure. Other emerging methods, such as hybrid residual CNN–LSTM architectures (Ren et al., 2020), versatile large language model-based frameworks (Li et al., 2024; Liu et al., 2024; Ren et al., 2024) or dynamic graph models (Huang et al., 2023), have so far been applied only to traffic forecasting, not interpolation tasks.
@@ 3.1 Model
Network Design. We model our network such that each street segment is thought to be equipped with a traffic volume sensor measuring the number of vehicles passing by (target variable). We use “street segment” interchangeably with “sensor”, under the simplifying assumption of a one-to-one mapping. In practice, a street segment may have multiple sensors; however, since our framework treats all sensors on the same segment as recording the same traffic volume, extending our approach to handle multiple sensors per segment is straightforward. In our experiments, these sensors are simulated based on fully observed proxy datasets (taxi and Strava counts), allowing us to emulate different levels of sensing coverage. The nodes in our graph represent the $n$ sensor locations in the training data (Figure 3). The edges and their weights are determined by the adjacency matrix $W\in\mathbb{R}^{n\times n}$. The adjacency matrix can represent distance or similarity, resulting in a fully connected graph (lower left of the figure), or be binary, with edges only between direct neighbors (lower right).
@@
Prediction Task. The goal of our model is to interpolate the traffic volume for street segments that lack sensor data. For this, the model is provided with a) the target variables and node features of seen (i.e., sensor-equipped), and b) the node features of the unseen (i.e., not equipped with sensors) segments. All predictions are made over a fixed time window. The model predicts the target variable for unseen and seen nodes across the time window. Wu et al. (2021a) compute the loss across all nodes, aiming to generalize the message-passing mechanism. Instead, we want our model to specialize in interpolating new locations. Thus, we compute the loss and the error metrics only on the interpolated nodes, while accounting for missing data (mathematical definition in A). For example, in Figure 2(a), the model uses the node features and the target variable of segments 1 to 7 (seen), to predict the target variable for segments 8 and 9 (unseen) on which the error metrics are computed.
Node Features and Target Variables. While most interpolation models rely primarily on the target variable from available sensors, we also incorporate node-level features. Features are important in urban contexts, where local characteristics significantly impact traffic volume. For example, a busy shopping street and a quiet residential street may be adjacent but have very different traffic volumes, differences that the model can only learn if features are included. All continuous input features, including the sensor measurements used as input to the model, are min-max scaled, and categorical features are one-hot encoded. No weighting is applied to any input features. We incorporate $k$ node features in our model. The data is available across $p$ time steps, the tensor $X\in\mathbb{R}^{n\times p\times k}$ contains the training node feature inputs for all sensor locations, while the target variable is stored in the tensor $T\in\mathbb{R}^{n\times p\times 1}$.
Time Window Approach to Model Temporal Dependency. To capture temporal dependencies, we adopt the fixed-length time window strategy from Wu et al. (2021a). For a given start time $t$, the model interpolates traffic volumes for the period $t,t+h)={t,t+1,\dots,t+h-1}$, where $h$ is the time window size. This assumes temporal correlation of traffic patterns within each window. While the employed time window approach effectively captures recurring traffic patterns (e.g., daily or weekly cycles), it is less suited to dynamic or abrupt changes in traffic, such as those caused by accidents, construction or other outstanding events. This limitation could be mitigated by incorporating real-time contextual signals (e.g., event or incident data) or by adopting adaptive temporal mechanisms that dynamically reweight recent observations. Exploring such adaptations remains an important direction for future work.
@@
Training Through Masking Strategy. To train the model for interpolation, we adopt a masking strategy inspired by Wu et al. (2021a). We briefly summarize it here and provide full and technical details in B. Missing values are common in real-world sensor data and can be caused by malfunctions, maintenance, or construction activities. For instance, after removing outliers and accounting for gaps, 1.3% of the Strava data and 1.0% of the taxi data are missing. The masking strategy handles the missing data by replacing it with zero. During training, we simulate the interpolation process by constructing multiple training samples. A sample is a small training instance that includes a) a selected time window of length $h$, and b) a subset of nodes, randomly divided into supposedly seen and unseen nodes. Each sample mimics a localized spatio-temporal snapshot of the network, allowing the model to repeatedly learn the interpolation task from different perspectives. To mimic the real prediction setting, we mask the target variables of the supposedly unseen nodes by setting them to zero, just as we represent missing values in the test phase. The model is trained to predict these masked values using the target variables and node features of the supposedly seen and the node features of the supposedly unseen nodes. Loss is computed only on the masked nodes. Also different from Wu et al. (2021a), we ensure that time windows during training do not overlap, to make use of all training data and adopt a scheduled learning rate to improve convergence and stability.
Setting target variables to zero can be problematic because the model cannot distinguish between genuinely zero values, those set to zero due to masking, and those set to zero because they are missing. To resolve this ambiguity, and differing from the original masking strategy in Wu et al. (2021a), we include two binary indicators in each node’s feature vector to denote masking and missingness. These are treated as standard input features and are already accounted for in the feature vector dimensionality k (see B for a mathematical formulation).
@@
Test Data from Same Time as Training Data. The test data comes from the same time period as the training data (in both Figure 2(a) and 2(b)). We train the model on time steps $1$ through $p$. The test is then performed on segments 8 and 9 during the time period $1$ through $p$. Since the test time overlaps with training, the model has already seen parts of the input (i.e., $X_{t}^{s}$, $X_{t}^{u}$, and $T_{t}^{s}$ under consideration of truly missing values). While this setup focuses on spatial interpolation, it requires retraining the model if predictions are needed for a different time. This is particularly suited for retrospective urban planning applications, where historical traffic distributions inform infrastructure decisions. In contrast, an alternative setup could consist of the model generalizing to both unseen locations and future times, meaning that no test inputs are observed during training. This is valuable in applications requiring continuously updated traffic forecasts, such as dynamic traffic control. This is explained and parallel to the results in this study, evaluated in an ablation study in C.
@@ 3.2 Adjacency matrices
Graph Neural Networks work with graph-structured data and, therefore, greatly depend on how the graph is defined (Ye et al., 2022). Various approaches exist for constructing adjacency matrices, including fixed, dynamic, or attention-based methods (i.e., graph attention networks). Fixed adjacency matrices are widely used, as they provide an interpretable and consistent neighborhood structure, particularly useful when modeling traffic systems with relatively stable spatial layouts, and are computationally efficient. However, fixed adjacency matrices fail to capture evolving spatial dependencies. Alternative strategies, such as learning dynamic adjacency matrices during training (e.g., Zhang et al., 2025), may better capture evolving traffic conditions, but at a higher computational cost. Similarly, graph attention networks (e.g., Zheng et al., 2020) offer another alternative by learning neighbor relevance directly from data. These models can be computationally efficient at scale, making them promising for large graphs, though they come with added architectural complexity. Given these trade-offs, we focus on fixed adjacency matrices in this work. Specifically, we investigate three types of fixed adjacency matrices commonly used in traffic modeling: binary, distance-based, and similarity-based. These were selected for their ease of construction and theoretical grounding in spatial and functional relationships between traffic sensors.
Binary adjacency matrices capture the local topology of the urban street network by connecting nodes (i.e., street segments) that are directly linked in space. They are widely used in the traffic domain due to their simplicity and ability to preserve physical road connectivity (Guo et al., 2019; Ye et al., 2022; Zhao et al., 2020). A binary adjacency matrix is defined using street segments $v_{i}$ and $v_{j}$ as:
$w_{ij}=\begin{cases}1,&\text{if }v_{i}\text{ and }v_{j}\text{ meet at an intersection},\\ 0,&\text{otherwise}.\end{cases}$ (1)
@@
[…]